OpenStudy (anonymous):

Which of the following is not an example of using the distributive property to multiply 12 times 22?

OpenStudy (texaschic101):

kinda need to see " which of the following " :)

OpenStudy (anonymous):

12(20 + 2) 22(10 + 2) 10(2 • 22)

OpenStudy (texaschic101):

10(2 * 22) is incorrect

OpenStudy (anonymous):

what about 22(10+20) or 12(20+2)

OpenStudy (texaschic101):

12(20 + 2) = 12 * 22 22(10 + 2) = 22 * 12 These are correct

OpenStudy (anonymous):

thanks :)

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

13(10 + 2) could be used to simplify which of the following problems? 130(26) 13(12) 13(20)

OpenStudy (texaschic101):

13(12)